### v2

# Retrieve Only When It Needs: Adaptive Retrieval Augmentation for Hallucination Mitigation in Large Language Models

Abstract
Hallucinations present a significant challenge for large language models (LLMs). The utilization of parametric knowledge in generating factual content is constrained by the limited knowledge of LLMs, potentially resulting in internal hallucinations. While incorporating external information can help fill knowledge gaps, it also introduces the risk of irrelevant information, thereby increasing the likelihood of external hallucinations. To balance the use of parametric knowledge within LLMs and external information, in this study, we present Rowen, a novel framework that enhances LLMs with an adaptive retrieval augmentation process tailored to address hallucinated outputs. Rowen introduces a consistency-based hallucination detection module, which assesses the model’s uncertainty regarding the input query by evaluating the semantic inconsistencies in various responses generated across different languages or models. When high uncertainties in the responses are detected, Rowen activates the retrieval of external information to rectify the model outputs. Through comprehensive empirical experiments, we demonstrate that Rowen surpasses the current state-of-the-art in both detecting and mitigating hallucinated content within the outputs of LLMs111Our code is available at https://github.com/dhx20150812/Rowen..
Retrieve Only When It Needs: Adaptive Retrieval Augmentation for Hallucination Mitigation in Large Language Models
Hanxing Ding1,2 Liang Pang1††thanks: Corresponding author Zihao Wei1,2 Huawei Shen1,2 Xueqi Cheng1,2 1Institute of Computing Technology, Chinese Academy of Sciences 2 University of Chinese Academy of Sciences {dinghanxing18s, pangliang, weizihao22z, shenhuawei, cxq}@ict.ac.cn
1 Introduction
In recent years, large language models (LLMs) have demonstrated impressive abilities in natural language understanding Hendrycks et al. (2021); Huang et al. (2023c), generation Touvron et al. (2023); Taori et al. (2023), and reasoning Zhang et al. (2023e); Wang et al. (2023a); Chu et al. (2023). Despite their successes, it has been widely observed that even state-of-the-art LLMs often generate factually incorrect or nonsensical outputs, referred to as hallucinations Ji et al. (2023a); Zhang et al. (2023d, b). These unreliable outputs pose significant risks in practical deployments of LLMs.
Efforts to enhance the factual accuracy of LLM outputs have been substantial. These studies often utilize LLMs’ extensive parametric knowledge and advanced logical reasoning capabilities. They employ approaches like self-reflection Wang et al. (2023b); Ji et al. (2023b); Madaan et al. (2023); Dhuliawala et al. (2023) or collaborative refinements involving interactions among multiple models Cohen et al. (2023); Du et al. (2023), aiming to enhance logical coherence in refined content. Despite their effectiveness, these self-improvement methods may be limited by LLMs’ knowledge boundaries Ren et al. (2023); Li et al. (2023a) or may not fully exploit parametric knowledge Huang et al. (2023a), leading to what we term internal hallucination, illustrated in Figure 1.
Alongside these self-improvement strategies, Retrieval-Augmented Generation (RAG) Chern et al. (2023) serves as a complementary method to overcome knowledge limitations. RAG employs a retrieve-then-read pipeline Karpukhin et al. (2020); Lewis et al. (2020), integrating relevant documents from external knowledge sources into the LLMs’ generation process Huang et al. (2023a); Ye et al. (2023); Varshney et al. (2023); Chern et al. (2023); Li et al. (2023b). However, as depicted in Figure 1, RAG methods are susceptible to external hallucination when irrelevant evidence is incorporated, potentially leading to cumulative errors and compromising output accuracy Li et al. (2023a); Shi et al. (2023).
Drawing inspiration from the latest neuroscience research Poskanzer and Aly (2023), which reveals how the human brain dynamically switches between internal thoughts and external sensations, we introduce a novel method, termed Rowen (Retrieve only when it needs). Rowen involves an innovative consistency-based uncertainty estimation module, which perturbs semantically equivalent questions and then evaluates the semantic inconsistency of various responses across diverse languages and models when subjected to these perturbed queries. To mitigate internal hallucinations, we trigger a retrieval process to fetch relevant information when Rowen detects inconsistencies in LLMs’ responses, indicating internal reasoning failures. This helps LLMs refine their reasoning chains and rectify potential hallucinations. To reduce external hallucinations, Rowen minimizes the risk of incorporating erroneous information by optimizing retrieval phases. If the perturbed answers convey consistent content, suggesting that LLMs are capable of generating the correct answer themselves, we directly adopt the original answer produced by internal reasoning. This method integrates parametric knowledge within LLMs with retrieved sources, ensuring a balanced integration of internal reasoning and external evidence to effectively mitigate hallucinations.
We evaluate the effectiveness of Rowen on the TruthfulQA dataset Lin et al. (2022) and the StrategyQA dataset Geva et al. (2021). Remarkably, our approach excels on the TruthfulQA dataset, yielding an impressive GPT-Judge score of 59.34%, marking a substantial improvement over the state-of-the-art (SOTA) baseline by a significant margin (+16.74%). Similarly, on the StrategyQA dataset, our approach achieves an accuracy of 75.60%, surpassing existing self-improvement and RAG-based baselines with notable superiority. These results comprehensively underscore the powerful capability of Rowen in mitigating hallucinated outputs of LLMs. Furthermore, our adaptive retrieval strategy significantly reduces unnecessary retrievals, thereby enhancing the efficiency of RAG systems.
2 Related Works
Mitigating hallucinations in the inference time could be a cost-effective and controllable way. A line of research harnesses the extensive parametric knowledge and robust logical reasoning capabilities of LLMs to ensure logical consistency either through self-reflection within a single model Wang et al. (2023b); Ji et al. (2023b); Xu et al. (2023); Madaan et al. (2023); Dhuliawala et al. (2023) or through collaborative refinements or debates involving multiple models Cohen et al. (2023); Du et al. (2023). Despite their strengths, LLMs are sometimes constrained by their knowledge boundaries and the complexity of the reasoning chain, resulting in occasional inaccuracies Ren et al. (2023); Li et al. (2023a); Mallen et al. (2023) termed internal hallucination. To address this knowledge gap, retrieval-augmented generation methods leverage external knowledge as supplementary evidence to aid LLMs in providing accurate responses Huang et al. (2023a); Ye et al. (2023); Niu et al. (2012); Xu et al. (2024); Varshney et al. (2023); Chern et al. (2023); Li et al. (2023b). However, these approaches, while effective, occasionally encounter the challenge of error accumulation, where irrelevant evidence may seep into the generation process, leading to incorrect responses Li et al. (2023a); Shi et al. (2023), a phenomenon referred to as external hallucination. Our work only performs retrieval augmentation when hallucinations are detected, thereby maximizing the utilization of both the parametric knowledge and externally retrieved information.
There are also some adaptive retrieval methods that assess the difficulty of questions or the confidence in responses to decide whether to retrieve documents Jiang et al. (2023); Mallen et al. (2023); Asai et al. (2023); Jeong et al. (2024). Due to space limitations, we discuss these works in Appendix C.
3 Methodology
Our objective is to enhance the factuality of LLM responses by integrating parametric and external knowledge. We propose a framework called Rowen (Retrieve only when needed). Initially, we leverage LLMs’ Chain-of-Thought (CoT) reasoning to generate an initial response (§ 3.1). To mitigate internal hallucinations, Rowen employs a consistency-based hallucination detection module that assesses the reliability of the initial response (§ 3.2). If high uncertainties are found, the initial answer is refined using external information via retrieval augmentation (§ 3.3), resulting in the final response. Otherwise, the initial response is considered the final output. For external hallucinations, Rowen resorts to external knowledge only when high uncertainties are found, ensuring that the final answer is both accurate and reliable.
3.1 Stage 1: Generating Initial Answer
To maximize the exploitation of the parametric knowledge in LLMs, we initially employ their Chain-of-Thought (CoT) reasoning to generate a preliminary response. This process involves: critically assessing the validity of the information in the input query and prioritizing accuracy and fact-checking for answer elaboration, detailed in Appendix F Table 9. After generating the CoT thought , we ask the model to provide a concise answer for the input query . The answer will serve as the final response to the input query if our detection module ensures it is free from hallucinations.
3.2 Stage 2: Deciding Whether to Retrieve
To decide when to retrieve, we leverage model uncertainty, which refers to the confidence level of model outputs and serves as a crucial indicator for deciding when to trust LLMs Zhang et al. (2023d). Unfortunately, current consistency-based methods fail when LLMs provide consistent yet incorrect answers across different perturbations Zhang et al. (2023a). This issue may arise because these methods focus exclusively on semantic coherence within a single language or model.
To tackle this issue, we propose novel cross-language and cross-model detection modules that assess semantic consistency among responses for the same question across different languages or models. If inconsistencies are detected in these responses, we flag them as potentially inaccurate and invoke a retrieval process.
3.2.1 Diversifying Question Verbalizations
To facilitate subsequent consistency-based hallucination detection, we begin by leveraging advancements in LLM prompting to generate semantically equivalent perturbations. Initially, we start with an input and instruct the model to provide a set of semantically equivalent questions . We use the prompt: "For the question [ORIGINAL QUESTION], please provide k semantically equivalent questions" with a high decoding temperature to generate diverse perturbed expressions.
After obtaining the diverse verbalized questions, we prompt the LM to generate its candidate responses according to the questions. We employ a greedy decoding strategy to avoid unpredictable randomness of the LM as much as possible.
where is the length of the generated semantically equivalent questions .
3.2.2 Cross-Language Consistency-based Detection
Recent finding reveals that multilingual models do not always learn knowledge that is fully aligned across different languages Gao et al. (2024). Building on this insight, we propose that this inconsistency can be strategically used as an indicator for the necessity of information retrieval. When language models provide inconsistent answers to the same query in different languages, it signals a lack of certainty about the relevant knowledge. In such cases, retrieving external knowledge is crucial for accurate responses.
To achieve cross-language detection, we first incorporate a cross-language consistency check (Rowen-CL) to evaluate the semantic consistency of responses to the same question across different languages. We introduce language-level perturbations by asking the model to translate the source-language questions into corresponding paraphrased questions in the target language. The model is then instructed to generate corresponding answers to each question in the target language.
To captures language-level cross-checking uncertainties, we utilize the generated questions and answers from all previous stages to calculate a numerical consistency score to calculate a numerical consistency score. Let denote a semantic equivalence checking operator that takes two QA pairs as inputs and returns "True" if they are semantically equivalent, and "False" otherwise. We then map the response to a numerical semantic equivalent score: {"True" 1.0, "False" 0.0}. We leverage the model and utilize the prompt provided in Appendix F Table 10 to implement the cross-language checking operator and calculate the cross-language consistency score as:
where and denote the QA pairs in the source language and target language, respectively.
3.2.3 Cross-Model Consistency-based Detection
Besides language-level cross-checking, we also introduce a cross-model detection module (Rowen-CM) to evaluate the semantic consistency of responses to the same question across different models. We adopt an additional verifier LM for model-level cross-checking and instruct the verifier LM to provide answers for each semantically equivalent question:
3.2.4 Hybrid Consistency-based Detection
The different variants of Rowen capture various aspects of uncertainty in the original response, complementing each other effectively. We propose integrating the cross-language and cross-model consistency scores to create a unified hybrid consistency score:
where is a weight factor for the cross-model consistency score.
3.3 Stage 3: Retrieval Augmented Generation
If the consistency score falls below a threshold, it indicates possible hallucinated content in the original response . We then introduce a retrieval-augmented generation procedure.
Searching Relevant Knowledge
To help the LM correct errors, we search for supporting evidence from external sources like online webpages. We first ask the model to generate search queries for each paraphrased question in , following the prompt in Appendix F Table 12. These queries are input into the online search engine to retrieve relevant knowledge, denoted as , used for correcting factual errors in .
Repairing Hallucinated Contents
With the retrieved evidence , the model reviews the original thought process and initial answer . The aim is to identify and correct inaccuracies, producing the refined answer :
Appendix F Table 13 shows the prompt for repairing the original answer. The corrected answer serves as the final response to question .
4 Experimental Setup
4.1 Datasets and Evaluation Metrics
TruthfulQA
We use the TruthfulQA dataset Lin et al. (2022) to evaluate the ability of LLMs in generating truthful responses Zhang et al. (2023c); Kadavath et al. (2022). In our study, we focus on the generation task in TruthfulQA. Therefore, to evaluate the factuality of responses from LLMs, we calculate the GPT-Judge score, obtained by fine-tuning the babbage-002 model using the original fine-tuning data from their official repository222https://github.com/sylinrl/TruthfulQA. We also report the BLEU and Rouge-L scores to evaluate the lexical overlap between generated responses and ground-truth references.
StrategyQA
The StrategyQA dataset Geva et al. (2021) comprises crowdsourced yes / no questions that require multi-step reasoning for accurate answers. We follow previous work Jiang et al. (2023) to randomly sample 500 examples due to the cost consideration of running experiments. We also follow the settings of Wei et al. (2022) to generate both the reasoning process as well as the final answer. We present the exact-match accuracy of the generated yes / no answers compared to the gold-standard answers.
| Models | TruthfulQA | StrategyQA | ||
| GPT-Judge | BLEU | Rouge-L | Accuracy | |
| Vanilla LLMs | ||||
| ChatGPT (gpt-3.5-turbo) | 47.92 | 10.17 | 31.31 | 61.40 |
| Self-improvement Methods | ||||
| CoVe Dhuliawala et al. (2023) | 48.01 | 12.81 | 26.52 | 61.40 |
| Multi-agent Debate Du et al. (2023) | 50.83 | 3.94 | 21.05 | 65.73 |
| Self-Reflection Ji et al. (2023b) | 42.99 | 3.86 | 18.18 | 62.40 |
| Retrieval-augmented Methods | ||||
| Factool Chern et al. (2023) | 34.50 | 1.34 | 12.22 | 67.20 |
| Detect-and-Mitigate Varshney et al. (2023) | 49.98 | 3.17 | 18.59 | 56.94 |
| Adaptive Retrieval Methods | ||||
| FLARE Jiang et al. (2023) | 45.04 | 11.59 | 26.83 | 61.19 |
| Adaptive-Retrieval Mallen et al. (2023) | 45.55 | 8.87 | 26.75 | 62.50 |
| Self-RAG Asai et al. (2023) | 40.36 | 4.36 | 21.28 | 58.40 |
| Adaptive-RAG Jeong et al. (2024) | 46.02 | 10.29 | 26.24 | 68.50 |
| LUQ Zhang et al. (2024) | 55.08 | 5.79 | 21.44 | 71.00 |
| Our Framework | ||||
| Rowen-CL | 57.39 | 7.60 | 24.16 | 74.00 |
| Rowen-CM | 56.29 | 6.85 | 22.36 | 72.40 |
| Rowen-Hybrid | 59.34 | 15.27 | 31.15 | 75.60 |
4.2 Baseline Methods
We consider the following methods as our baselines: (1) Vanilla LLMs, such as ChatGPT. (2) Self-improvement methods: CoVe Dhuliawala et al. (2023) generates verification questions to self-analyze potential errors, systematically addressing each question to refine the baseline response. Self-Reflection Ji et al. (2023b) presents an interactive self-reflection methodology that incorporates knowledge acquisition and answer generation. Multi-agent Debate Du et al. (2023) utilize multiple LM agents to debate their individual responses over multiple rounds to arrive at a common final answer. (3) Retrieval-augmented methods: Factool Chern et al. (2023) leverages various tools to gather evidence about the factuality of the generated content. Detect and Mitigate Varshney et al. (2023) actively detects hallucinations during generation by identifying potential hallucination through the logit output values of LLMs. (4) Adaptive retrieval methods: FLARE Jiang et al. (2023) adopts an active retrieval strategy that only retrieves when LLMs generate low probability tokens. Adaptive-Retrieval Mallen et al. (2023) only retrieves when necessary based on a pre-defined threshold for entity popularity. Self-RAG Asai et al. (2023) trains a single arbitrary LM that adaptively retrieves passages on-demand. Adaptive-RAG Jeong et al. (2024) trains a query-complexity classifier to decide when to retrieve based on question complexity. LUQ Zhang et al. (2024) is a sampling-based uncertainty quantification for long text.
4.3 Implementation Details
In our experiments, we validate Rowen using the ChatGPT language model. We re-implement all baselines, except Self-RAG, using ChatGPT to ensure a fair comparison. For semantic perturbations, we configure the temperature to 1.0 to generate diverse perturbed expressions. Otherwise, the temperature is set to 0.0 to obtain high-quality deterministic outputs. Considering the diversity of expressions and the latency in generating perturbations, we produce semantically equivalent questions. For the cross-language detection module, English serves as the source language while Chinese is employed as the target language. For the cross-model detection module, we adopt Qwen-Max-0428333https://qwenlm.github.io/blog/qwen-max-0428, a large instruction-tuned model for chat service, as the verifier LM. Following Chern et al. (2023), we utilize the Google Search API offered by Serper444https://serper.dev/ to search the top pages and extract the most pertinent search snippets from the API’s response. We also conduct extensive experiments to study the impact of various hyper-parameters in Appendix D.
5 Experimental Results
5.1 Main Results on Hallucination Mitigation
We evaluate the effectiveness of Rowen on the TruthfulQA and StrategyQA datasets. Table 1 presents the overall performance of Rowen compared to several strong baselines. Rowen demonstrates superior performance on both datasets, with higher GPT-Judge score and accuracy, indicating the effectiveness of our proposed method.
| Models | TruthfulQA | StrategyQA | ||||
| GPT-Judge | BLEU | Rouge-L | Ratio(%) | Accuracy | Ratio(%) | |
| LLaMa2-7B | ||||||
| Average Token Probability | 50.19 | 12.51 | 31.11 | 46.5 | 71.50 | 23.0 |
| Average Token Entropy | 52.22 | 9.18 | 28.37 | 59.0 | 72.00 | 26.5 |
| SelfCheckGPT | ||||||
| w/ BERTScore | 51.38 | 8.43 | 26.85 | 27.2 | 67.50 | 21.0 |
| w/ MQAG | 52.76 | 7.69 | 26.92 | 54.4 | 69.00 | 34.0 |
| w/ Ngram | 52.41 | 5.01 | 21.60 | 34.9 | 66.50 | 32.0 |
| Combination | 53.10 | 6.69 | 24.04 | 51.3 | 69.50 | 30.0 |
| Consistency | ||||||
| SAC3-Q | 51.02 | 7.90 | 28.00 | 24.5 | 65.50 | 24.0 |
| SAC3-all | 52.22 | 9.37 | 29.56 | 20.8 | 67.00 | 24.5 |
| Rowen-Hybrid | 59.34 | 15.27 | 31.15 | 23.0 | 75.6 | 20.0 |
Vanilla ChatGPT shows a certain level of accuracy in answering factual questions, achieving scores of 47.92% and 61.40% on the two datasets, respectively. While self-improvement methods perform better than the vanilla LM on both datasets, they are still limited by their knowledge boundaries and suffer from internal hallucinations. RAG methods demonstrate relatively better performance compared to self-improvement methods, highlighting the benefits of integrating external knowledge. However, Factool falls short on the TruthfulQA dataset, and the Detect-and-Mitigate method underperforms on the StrategyQA dataset. This may be attributed to error accumulation caused by unnecessary retrieval (external hallucinations).
We also conduct additional experiments to compare with four adaptive retrieval methods. Notably, Adaptive-Retrieval faces challenges on the TruthfulQA dataset due to some questions lacking explicit entities, causing it to struggle in deciding when to retrieve based on entity popularity, leading to poor performance. Besides, Self-RAG’s effectiveness is hindered by the capabilities of the LLaMa model, resulting in inferior performance.
Compared to the aforementioned baselines, Rowen demonstrates significant performance gains on both datasets. Both Rowen-CL and Rowen-CM exhibit excellent hallucination mitigation capabilities, even when compared to strong adaptive retrieval methods. Specifically, Rowen-Hybrid achieves a GPT-Judge score of 59.34% on the TruthfulQA dataset, surpassing the strongest baseline by 16.74%. Additionally, Rowen-Hybrid attains an accuracy of 75.60% on the StrategyQA dataset, significantly outperforming existing baselines. These results underscore Rowen’s ability to effectively leverage parametric knowledge and external information for maximum advantage.
5.2 Effect of Detection Module
To validate the superiority of our hallucination detection module, we compare its performance in adaptive retrieval scenarios for hallucination mitigation with strong detection methods: (1) average token-level probability / entropy that utilizes the probabilities / entropies of tokens generated by a proxy LM (e.g., LLaMa2-7B) as a metric to measure hallucination. (2) SelfCheckGPT Manakul et al. (2023) that measures information consistency between the different responses to determine hallucinations. (3) Consistency-based method Zhang et al. (2023a), SAC3, that evaluates semantic-aware cross-check consistency, building upon the foundation of self-consistency principles.
Based on the results in Table 2, it is evident that logits-based methods perform moderately well in detecting hallucinations, especially on the StrategyQA dataset. Specifically, the entropy-based estimation exhibits superior performance. Besides, the combination of different variants in SelfCheckGPT leads to a slight performance improvement. Additionally, it is worth noting that SAC3 achieve competitive performance on detecting potential factual errors accurately.
Finally, we observe that Rowen significantly outperforms these strong hallucination detection baselines, especially the monolingual detection method SAC3. We also notice that Rowen achieves notable hallucination mitigation with a minimal number of retrieval calls. This underlines the superior efficiency of our hallucination detection module.
5.3 Scalability of Rowen
Scalability to Open-Source LLMs
In addition to ChatGPT, we also assess Rowen’s effectiveness when employing open-source language models: Qwen1.5-14B-Chat, Qwen2-7B-Instruct, and Qwen2-72B-Instruct555https://huggingface.co/Qwen. These models are chosen for the cross-language detection model due to their strong capabilities in following Chinese instructions, a critical feature for effective cross-language detection. The Llama-series models are not considered due to their weaker performance in generating Chinese responses. Instead, the Qwen-series models have demonstrated state-of-the-art performance in Chinese, which aligns with our research goals. For verifier LM , we choose to use Llama-3-8B-Instruct666https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct.
Figure 3 shows the results of three variants of Rowen on three open-source models on TruthfulQA dataset. Our Rowen methods achieve strong performance when applied to open-source LLMs, nearly matching the baseline results on ChatGPT. Specifically, the Multi-agent Debate baseline scores 50.83, while the Detect-and-Mitigate baseline scores 49.98. These findings further prove the effectiveness and scalability of our proposed Rowen method within open-source LLMs.
Scalability to Other Datasets
We assess the Rowen model’s scalability by examining its performance on datasets with answers of intermediate length, namely TriviaQA and Natural Questions, given the diverse answer lengths in TruthfulQA (long answers) and StrategyQA (binary responses). We conduct comparisons against two strong adaptive retrieval baselines, reporting the metrics of EM (Exact Match) and F1 score on their dev sets. The experimental results are shown in Table 3.
In short, our Rowen model, with its different variants, consistently outperforms the baselines on both datasets, suggesting a notably more robust capability in handling intermediate-length answers. Particularly, Rowen-Hybrid emerges as the most effective, achieving the highest scores across both metrics and datasets. This denotes a significant enhancement over baseline methods.
| Methods | NQ | TriviaQA | ||
| EM | F1 | EM | F1 | |
| FLARE | 32.50 | 43.91 | 59.00 | 68.34 |
| Adaptive-RAG | 35.04 | 48.44 | 58.00 | 68.97 |
| Rowen-CL | 38.08 | 55.81 | 66.50 | 76.66 |
| Rowen-CM | 37.36 | 53.27 | 65.00 | 74.98 |
| Rowen-Hybrid | 39.98 | 57.31 | 69.04 | 78.46 |
5.4 Analysis of Computation Cost for Rowen
Efficiency Analysis of LLM Calls
To evaluate the efficiency of LLM calls in Rowen, we compare various methods—including Self-Reflection, Multi-agent Debate and three different Rowen variants—by analyzing the number of API calls required to answer a single question and their effectiveness in mitigating hallucinations. From the results shown in Table 4, we find that Rowen variants, particularly Rowen-CL and Rowen-CM, achieve significantly better hallucination mitigation compared to the baseline methods, while using a similar number of API call. To further improve efficiency, We design two useful speedup strategies to improve the efficiency of Rowen and describe them in Appendix E.
| Methods | TruthfulQA | StrategyQA | ||
| GPT-Judge | # Calls | Accuracy | # Calls | |
| Self-Reflection | 42.99 | 6 | 62.40 | 5 |
| Multi-agent Deb. | 50.83 | 6 | 65.73 | 6 |
| Rowen-CL | 57.39 | 6 | 74.00 | 5 |
| Rowen-CM | 56.29 | 5 | 72.40 | 4 |
| Rowen-Hybrid | 59.34 | 8 | 75.60 | 6 |
Efficiency Analysis of Retrieval Calls
To verify the retrieval efficiency of Rowen, in Table 5, we compare the average number of retrieval calls made by Factool, Detect-and-Mitigate, FLARE, and Rowen-Hybrid to answer a question across two datasets. Rowen-Hybrid, which retrieves information only for uncertain responses, excels in both retrieval efficiency and factual accuracy, showcasing its superiority over other RAG methods.
| Methods | # Num of Retrieval Calls | |
| TruthfulQA | StrategyQA | |
| Factool | 12.5 | 11.6 |
| Detect-and-Mit. | 7.2 | 5.5 |
| FLARE | 2.1 | 3.9 |
| Adaptive-RAG | 0.9 | 1.4 |
| Rowen-Hybrid | 1.5 | 0.5 |
5.5 Quantitative Analysis of Hallucinations
To verify whether Rowen effectively reduces internal and external hallucinations, we present a comparative analysis of the prevalence of both types of hallucinations across three methods—Factool, Detect-and-Mitigate, and Rowen-Hybrid. Internal hallucinations refer to the generation of wrong responses using only the parameterized knowledge of LLMs, while external hallucinations refer to the generation of incorrect responses using noisy documents introduced after retrieval. Figure 4 shows that Factool and Detect-and-Mitigate are significantly prone to both types of hallucinations. Both baseline methods have high levels of internal and external hallucinations, struggling with internal coherence and external fact alignment. In contrast, Rowen-Hybrid effectively reduces external hallucinations by timely integrating external knowledge, thereby avoiding unnecessary information retrieval and mitigating potential errors.
6 Conclusion and Future Work
In this paper, we introduce Rowen, a novel method designed to actively identify potential hallucinated responses in LLMs and correcting factual errors through an adaptive retrieval process. Our experiments on four datasets demonstrate the effectiveness of our approach. We also conduct detailed analytical experiments to underscore the impact of each module and the choices of hyper-parameters. Moving forward, we plan to expand our work by exploring how to effectively utilize retrieved evidence, even when irrelevant documents are integrated, and devise an effective method to minimize the influence of knowledge conflicts between internal knowledge and retrieved documents.
Limitations
Our method has several drawbacks. Firstly, it focuses on post-hoc correction, limiting its ability to detect hallucination only after generating a complete sentence rather than in real-time during the generation of contexts. Secondly, for ensuring accurate answers, our approach requires multiple requests to the API interface. This involves generating the CoT answer, perturbing and responding to the input, correcting the original answer, and so on, leading to increased latency. To address this, we can expedite the entire process by parallelizing API calls for both the OpenAI API and Google Search API. Additionally, reducing the number of perturbations can strike a better balance between detection precision and inference speed. Thirdly, our approach neglects to analyze and utilize the factual nature of the retrieved evidence, which we leave for future work.
Ethics Statement
This study delves into the detection and mitigation of hallucinations in language models (LMs), presenting broader implications for the realm of natural language processing (NLP) and addressing ethical concerns related to trust and reliability. Our method has the potential to enhance the accuracy and dependability of LMs, minimizing the risks associated with misinformation and biased outputs. Moreover, it foster accountability and trust in AI systems, contributing to the ongoing development of more reliable language models.
Appendix A Kernel Density Estimation Visualization
In our work, we hypothesize that lower consistency levels correlate with increased uncertainty in the model, consequently leading to a higher probability of inaccurate predictions. To verify this hypothesis, we analyze the consistency distribution through kernel density estimation plots on the TruthfulQA dataset. As illustrated in Figure 5, in the lower range of consistency check scores (left-skewed), the proportion of incorrect predictions is higher. This indicates that low consistency check scores are often associated with incorrect model predictions. On the contrary, in the higher range of consistency check scores (right-skewed), the proportion of correct predictions is higher. This indicates that high consistency check scores are usually associated with correct model predictions. There is some overlap between the distributions of different labels, but the overall trend shows that the higher the consistency check score, the higher the likelihood of correct predictions. These findings suggest that consistency check scores can be an effective indicator of prediction accuracy: the higher the score, the more accurate the model prediction is likely to be. These findings provide empirical support for our hypothesis. Additionally, we observe that the suitable threshold points for distinguishing correct from incorrect answers differ between Rowen-CL and Rowen-CM, with the optimal threshold for Rowen-CM being higher. This observation aligns with the conclusion in § D.
Appendix B Related Work on Exploring Uncertainty for Hallucination Detection
Uncertainty refers to the confidence level of the model outputs, and it serves an important indicator for identifying and eliminating hallucinations, so it can assist users in determining when to trust LLMs. In general, methods for exploring uncertainty for hallucination detection can be categorized into three types: (1) Logit-based estimation relies on accessing the model’s logits to calculate token-level probabilities or entropy, which are used to measure uncertainty Huang et al. (2023b); Varshney et al. (2023). However, this approach can pose challenges for black-box closed-source models. (2) Verbalized-based estimation involves prompting language models to express their uncertainty using specific prompts Xiong et al. (2023); Tian et al. (2023); Agrawal et al. (2023). However, these methods tend to display a high degree of overconfidence when expressing their confidence Xiong et al. (2023); Tao et al. (2024). (3) To overcome these limitations, consistency-based estimations are adopted to measure the consistency score among multiple responses provided by the model for a given question Manakul et al. (2023); Xiong et al. (2023); Wang et al. (2023b); Zhao et al. (2023). The underlying assumption suggests that when language models struggle with indecision and fabricate facts, they tend to provide logically inconsistent responses to identical questions. In this work, we propose that cross-language and cross-model consistency can offer highly sensitive signals for identifying hallucinations. Therefore, we utilize cross-language and cross-model detection modules that cross-check answers to the same question across different languages or models. This cross-checking paradigm serves as a powerful mechanism to identify hallucinations in LLMs.
Appendix C Discussions
There is a line of recent studies that concentrate on mitigating hallucinations with adaptive retrieval Jiang et al. (2023); Mallen et al. (2023); Asai et al. (2023); Jeong et al. (2024). Our work diverges from existing methodologies by: (1) Traditional self-consistency method Wang et al. (2023b) may falter when LLMs provide coherent yet incorrect responses to the same query, which is observed in SAC3. Our approach extends self-consistency uncertainty estimation to multilingual and multi-model settings, yielding superior hallucination mitigation performance compared to conventional methods (see Table 1). (2) FLARE Jiang et al. (2023) relies on the token probability of model outputs, a limitation for commercial closed-source models. In contrast, our method overcomes this constraint by assessing semantic similarity in generated content, which is suitable for white / black-box models. (3) Self-RAG Asai et al. (2023) necessitates intricate retraining processes and significant computational resources (4 Nvidia A100 GPUs and 80GB memory). In contrast, our method effectively mitigates hallucination without the need for retraining, underscoring its suitability for resource-constrained scenarios. (4) Adaptive-Retrieval Mallen et al. (2023) employs Wikipedia entity popularity for retrieval decisions, which proves ineffective for adversarial questions lacking clear entities in datasets like TruthfulQA. (5) Adaptive-RAG Jeong et al. (2024) dynamically decide whether to retrieve based on a pre-trained classifier, which is a smaller LM trained to predict the complexity level of incoming queries. Adaptive-RAG could be influenced by the accuracy of the classifier, given the absence of training datasets for query-complexity classifier, potentially leading to incorrect labeling of queries.
Our approach, examining linguistic diversity (cross-language) and model response consistency (cross-model), effectively identifies and mitigates hallucinations. This enhances model robustness and reliability, providing a solid foundation for applying LLMs. Our method is novel, practical, flexible, training-free, and broadly applicable, offering new insights and optimizing LLM research and application.
Appendix D Impact of Hyper-parameters in Rowen
Impact of Detection Threshold.
We experiment with various consistency thresholds to investigate their impact on hallucination mitigation effectiveness. As shown in Figure 6, we observe significant performance improvement when the detection threshold increases from 0.2 to 0.6, as this identifies more hallucinations and enhances factual accuracy by retrieving more external information. However, further increasing the threshold introduces more noise and irrelevant evidence, degrading performance. This suggests that an appropriate threshold is crucial for balancing detection accuracy and the retrieval process. Additionally, Rowen-CL and Rowen-CM have different optimal thresholds of 0.6 and 0.8, respectively, and the Rowen-Hybrid method demonstrates the most stable performance across varying thresholds.
Impact of Number of Perturbed Questions.
We investigate the impact of the number of perturbed questions of Rowen on the performance of hallucination mitigation. The results in Figure 7 reveal that the mitigation performance improves with the increase in the number of perturbations, particularly notable in the transition from 4 to 6 perturbations. This enhancement is attributed to the greater number of perturbations aiding in higher detection accuracy. However, further increasing the perturbation count beyond 6 yields only marginal improvements in performance (even no improvements). Thus, after considering the results, we determine that the optimal number of perturbations for achieving the best hallucination mitigation performance is 6.
Impact of Cross-Model Detection Weight.
We also examine the impact of different cross-model detection weight for Rowen-Hybrid. We ensure that for any given , the detection threshold is recalibrated to maintain optimal performance. Figure 8 shows that GPT-Judge scores significantly increase as the weight rises from 0.25 to 1.0, indicating that introducing cross-model checking can improve the accuracy of hallucination detection. However, when the weight factor exceeds 1.0, performance declines, with scores dropping to around 58 at a weight of 1.5. This suggests that while a moderate weight enhances factuality, excessively a high value allows the cross-model detection module to exert too much influence, reducing effectiveness.
Impact of the Choice of Target Language
To investigate how different language pairs affect the effectiveness of hallucination mitigation for Rowen-CL, we conduct experiments on the TruthfulQA dataset with various target languages. The results of these experiments are presented in Table 6. It is observed that the combination of English as the source language and German as the target language yields the least favorable results. This may be attributed to their shared Germanic language family roots, which results in numerous linguistic similarities and overlapping cultural references. Conversely, as the cultural divergence between the source and target languages widens, we witness an enhancement in the performance of hallucination mitigation. This trend substantiates the hypothesis that cultural disparities between languages play a pivotal role in identifying hallucinations and bolstering the factuality of the generated responses.
| Language Pair | Discrepancy | GPT-Judge |
| English - German | Low | 50.55 |
| English - French | Medium | 52.26 |
| English - Chinese | High | 57.39 |
Impact of the Choice of Verifier LM
To investigate the impact of different verifier LMs on the efficacy of hallucination mitigation for Rowen-CM, we conducted experiments using various verifier LMs on the TruthfulQA dataset. The experimental results presented in Table 7 demonstrate a significant improvement in hallucination mitigation as the capability of the verifier LM increases. Specifically, Qwen-Max exhibited the highest effectiveness in this task, achieving an efficacy of 56.29%, while Qwen-Turbo and Qwen-Plus achieved 52.50% and 54.09%, respectively. These findings underscore the critical role of verifier LM selection in enhancing the effectiveness of hallucination mitigation.
| Verifier LM | Capability | GPT-Judge |
| Qwen-Turbo | Low | 52.50 |
| Qwen-Plus | Medium | 54.09 |
| Qwen-Max | High | 56.29 |
Impact of Perturbation Objects
In evaluating cross-language and cross-model perturbations, as well as calculating subsequent consistency scores, applying CL and CM perturbations to the original question appears to be a viable approach. However, recent studies suggest that merely perturbing questions might lead to responses that are consistent but not necessarily accurate Zhang et al. (2023a). This implies that question-level consistency alone does not suffice for assessing factual correctness, as these methods predominantly emphasize semantic coherence.
To validate this, we conduct preliminary experiments by applying CL and CM perturbations to the original question (question-level consistency checking) and compare the results with those of Rowen. The results, as shown in Table 8, demonstrate that Rowen consistently outperforms the approach of applying perturbations to the original question. This reinforces that Rowen remains the most effective method for ensuring accuracy.
| Object | Variants | TruthfulQA | StrategyQA |
| GPT-Judge | Accuracy | ||
| Input | CL | 55.34 | 72.00 |
| CM | 55.85 | 72.00 | |
| Perturbations | CL | 57.39 | 74.00 |
| CM | 56.29 | 72.40 |
Appendix E Strategies to improve efficiency
We have implemented two helpful strategies to improve the efficiency without compromising the performance by:
(1) Our cross-language detection and cross-model detection modules can operate in parallel. This significantly reduces latency, as it allows for the simultaneous evaluation of semantic consistency across different languages and models.
(2) We have optimized the generation process for responding to multiple semantically equivalent perturbed questions by leveraging parallel API calls. This approach expedites the retrieval of diverse responses, which are crucial for our consistency checks.
By employing these parallel processing strategies, we ensure that the efficiency of our method is improved while preserving its effectiveness. We believe these optimizations adequately address the concerns about the method’s efficiency and demonstrate our commitment to developing a scalable and high-performing framework.
Appendix F Prompt Instruction Examples
In this section, we provide these instructions used in our experiments.
| When responding to my question, please first evaluate the validity of the information or the assumption underlying the question. Once you’ve established its truth or existence, then proceed to deliver a detailed explanation or answer. Prioritize accuracy and fact-checking before diving into elaboration or conjecture. |
| Given the question Q, and two potential answers: answer A in English and answer B in Chinese. Your task is to determine if the content and meaning of A and B are equivalent in the context of answering Q. Consider linguistic nuances, cultural variations, and the overall conveyance of information. Respond with a binary classification. If A and B are equivalent, output ’True.’, otherwise output ’False’ |
| Are the following two Question-Answer(QA) pairs semantically equivalent? Provide your best guess that it is correct (True or False). Given ONLY the guess (True or False), no other words or explanation. |
| Transform the following question into a query suitable for information retrieval and then use the query to find relevant evidence. Follow these steps for a short but comprehensive approach: |
| Original Question: |
| 1. Simplify the question into key terms and concepts. |
| 2. Remove any elements that are not essential for a search query. |
| 3. If necessary, rephrase the question to focus on the most critical aspects for retrieval. |
| You are an advanced AI with the ability to process and synthesize information efficiently. When provided with a question, you should consider the following aspects in your response: |
| 1. The specific question asked: |
| 2. The reasoning process: |
| 3. The short answer previously given related to the question: |
| 4. Evidences that have been retrieved: |
| Use this information to generate a concise, accurate, and relevant answer that reflects the latest understanding of the topic based on the input provided. |
| Your response should be clear, direct, and provide the most up-to-date information available while maintaining coherence with the previous discussion points. |

### v3

@@
-Retrieve Only When It Needs: Adaptive Retrieval Augmentation for Hallucination Mitigation in Large Language Models
-Abstract
+Rowen: Adaptive Retrieval-Augmented Generation for Hallucination Mitigation in LLMs
+Abstract.
 Hallucinations present a significant challenge for large language models (LLMs).
@@
 When high uncertainties in the responses are detected, Rowen activates the retrieval of external information to rectify the model outputs.
-Through comprehensive empirical experiments, we demonstrate that Rowen surpasses the current state-of-the-art in both detecting and mitigating hallucinated content within the outputs of LLMs111Our code is available at https://github.com/dhx20150812/Rowen..
-Retrieve Only When It Needs: Adaptive Retrieval Augmentation for Hallucination Mitigation in Large Language Models
-Hanxing Ding1,2 Liang Pang1††thanks: Corresponding author Zihao Wei1,2 Huawei Shen1,2 Xueqi Cheng1,2 1Institute of Computing Technology, Chinese Academy of Sciences 2 University of Chinese Academy of Sciences {dinghanxing18s, pangliang, weizihao22z, shenhuawei, cxq}@ict.ac.cn
-1 Introduction
-In recent years, large language models (LLMs) have demonstrated impressive abilities in natural language understanding Hendrycks et al.
-(2021); Huang et al.
-(2023c), generation Touvron et al.
-(2023); Taori et al.
-(2023), and reasoning Zhang et al.
-(2023e); Wang et al.
-(2023a); Chu et al.
-(2023).
-Despite their successes, it has been widely observed that even state-of-the-art LLMs often generate factually incorrect or nonsensical outputs, referred to as hallucinations Ji et al.
-(2023a); Zhang et al.
-(2023d, b).
+Through comprehensive empirical experiments, we demonstrate that Rowen surpasses the current state-of-the-art in both detecting and mitigating hallucinated content within the outputs of LLMs111Our Code: https://github.com/dhx20150812/Rowen..
+1. Introduction
+In recent years, large language models (LLMs) have demonstrated impressive abilities in natural language understanding (Hendrycks et al., 2021; Huang et al., 2023a), generation (Touvron et al., 2023; Taori et al., 2023), and reasoning (Zhang et al., 2023e; Wang et al., 2023a; Chu et al., 2023).
+Despite their successes, it has been widely observed that even state-of-the-art LLMs often generate factually incorrect or nonsensical outputs, referred to as hallucinations (Ji et al., 2023a; Zhang et al., 2023b, d).
 These unreliable outputs pose significant risks in practical deployments of LLMs.
@@
 These studies often utilize LLMs’ extensive parametric knowledge and advanced logical reasoning capabilities.
-They employ approaches like self-reflection Wang et al.
-(2023b); Ji et al.
-(2023b); Madaan et al.
-(2023); Dhuliawala et al.
-(2023) or collaborative refinements involving interactions among multiple models Cohen et al.
-(2023); Du et al.
-(2023), aiming to enhance logical coherence in refined content.
-Despite their effectiveness, these self-improvement methods may be limited by LLMs’ knowledge boundaries Ren et al.
-(2023); Li et al.
-(2023a) or may not fully exploit parametric knowledge Huang et al.
-(2023a), leading to what we term internal hallucination, illustrated in Figure 1.
-Alongside these self-improvement strategies, Retrieval-Augmented Generation (RAG) Chern et al.
-(2023) serves as a complementary method to overcome knowledge limitations.
-RAG employs a retrieve-then-read pipeline Karpukhin et al.
-(2020); Lewis et al.
-(2020), integrating relevant documents from external knowledge sources into the LLMs’ generation process Huang et al.
-(2023a); Ye et al.
-(2023); Varshney et al.
-(2023); Chern et al.
-(2023); Li et al.
-(2023b).
-However, as depicted in Figure 1, RAG methods are susceptible to external hallucination when irrelevant evidence is incorporated, potentially leading to cumulative errors and compromising output accuracy Li et al.
-(2023a); Shi et al.
-(2023).
-Drawing inspiration from the latest neuroscience research Poskanzer and Aly (2023), which reveals how the human brain dynamically switches between internal thoughts and external sensations, we introduce a novel method, termed Rowen (Retrieve only when it needs).
+They employ approaches like self-reflection (Wang et al., 2023b; Ji et al., 2023b; Madaan et al., 2023; Dhuliawala et al., 2023) or collaborative refinements involving interactions among multiple models (Cohen et al., 2023; Du et al., 2023), aiming to enhance logical coherence in refined content.
+Despite their effectiveness, these self-improvement methods may be limited by LLMs’ knowledge boundaries (Ren et al., 2023; Li et al., 2023b) or may not fully exploit parametric knowledge (Huang et al., 2023c), leading to what we term internal hallucination, illustrated in Figure 1.
+Alongside these self-improvement strategies, Retrieval-Augmented Generation (RAG) (Chern et al., 2023) serves as a complementary method to overcome knowledge limitations.
+RAG employs a retrieve-then-read pipeline (Karpukhin et al., 2020; Lewis et al., 2020), integrating relevant documents from external knowledge sources into the LLMs’ generation process (Huang et al., 2023c; Ye et al., 2023; Varshney et al., 2023; Chern et al., 2023; Li et al., 2023a).
+However, as depicted in Figure 1, RAG methods are susceptible to external hallucination when irrelevant evidence is incorporated, potentially leading to cumulative errors and compromising output accuracy (Li et al., 2023b; Shi et al., 2023).
+Drawing inspiration from the latest neuroscience research (Poskanzer and Aly, 2023), which reveals how the human brain dynamically switches between internal thoughts and external sensations, we introduce a novel method, termed Rowen (Retrieve only when it needs).
 Rowen involves an innovative consistency-based uncertainty estimation module, which perturbs semantically equivalent questions and then evaluates the semantic inconsistency of various responses across diverse languages and models when subjected to these perturbed queries.
@@
 This method integrates parametric knowledge within LLMs with retrieved sources, ensuring a balanced integration of internal reasoning and external evidence to effectively mitigate hallucinations.
-We evaluate the effectiveness of Rowen on the TruthfulQA dataset Lin et al.
-(2022) and the StrategyQA dataset Geva et al.
-(2021).
+We evaluate the effectiveness of Rowen on the TruthfulQA dataset (Lin et al., 2022) and the StrategyQA dataset (Geva et al., 2021).
 Remarkably, our approach excels on the TruthfulQA dataset, yielding an impressive GPT-Judge score of 59.34%, marking a substantial improvement over the state-of-the-art (SOTA) baseline by a significant margin (+16.74%).
 Similarly, on the StrategyQA dataset, our approach achieves an accuracy of 75.60%, surpassing existing self-improvement and RAG-based baselines with notable superiority.
-These results comprehensively underscore the powerful capability of Rowen in mitigating hallucinated outputs of LLMs.
+These results underscore the powerful capability of Rowen in mitigating hallucinated outputs of LLMs.
 Furthermore, our adaptive retrieval strategy significantly reduces unnecessary retrievals, thereby enhancing the efficiency of RAG systems.
-2 Related Works
+2. Related Works
+In this section, we discuss recent works on hallucination detection and mitigation, focusing on uncertainty estimation methods for detection and post-hoc correction for mitigation.
+2.1. Exploring Uncertainty for Hallucination Detection
+Uncertainty refers to the confidence level of the model outputs, and it serves an important indicator for identifying and eliminating hallucinations, so it can assist users in determining when to trust LLMs.
+In general, methods for exploring uncertainty for hallucination detection can be categorized into three types: (1) Logit-based estimation relies on accessing the model’s logits to calculate token-level probabilities or entropy, which are used to measure uncertainty (Manakul et al., 2023; Huang et al., 2023b; Varshney et al., 2023).
+However, this approach can pose challenges for black-box closed-source models.
+(2) Verbalized-based estimation involves prompting language models to express their uncertainty using specific prompts (Xiong et al., 2023; Tian et al., 2023; Agrawal et al., 2023).
+However, these methods tend to display a high degree of overconfidence when expressing their confidence (Xiong et al., 2023; Tao et al., 2024).
+(3) To overcome these limitations, consistency-based estimations are adopted to measure the consistency score among multiple responses provided by the model for a given question (Manakul et al., 2023; Xiong et al., 2023; Wang et al., 2023b; Zhao et al., 2023).
+The underlying assumption suggests that when language models struggle with indecision and fabricate facts, they tend to provide logically inconsistent responses to identical questions.
+In this work, we propose that cross-language and cross-model consistency can offer highly sensitive signals for identifying hallucinations.
+Therefore, we utilize cross-language and cross-model detection modules that cross-check answers to the same question across different languages or models.
+This cross-checking paradigm serves as a powerful mechanism to identify hallucinations in LLMs.
+2.2. Post-hoc Correction for Hallucination Mitigation
 Mitigating hallucinations in the inference time could be a cost-effective and controllable way.
-A line of research harnesses the extensive parametric knowledge and robust logical reasoning capabilities of LLMs to ensure logical consistency either through self-reflection within a single model Wang et al.
-(2023b); Ji et al.
-(2023b); Xu et al.
-(2023); Madaan et al.
-(2023); Dhuliawala et al.
-(2023) or through collaborative refinements or debates involving multiple models Cohen et al.
-(2023); Du et al.
-(2023).
-Despite their strengths, LLMs are sometimes constrained by their knowledge boundaries and the complexity of the reasoning chain, resulting in occasional inaccuracies Ren et al.
-(2023); Li et al.
-(2023a); Mallen et al.
-(2023) termed internal hallucination.
-To address this knowledge gap, retrieval-augmented generation methods leverage external knowledge as supplementary evidence to aid LLMs in providing accurate responses Huang et al.
-(2023a); Ye et al.
-(2023); Niu et al.
-(2012); Xu et al.
-(2024); Varshney et al.
-(2023); Chern et al.
-(2023); Li et al.
-(2023b).
-However, these approaches, while effective, occasionally encounter the challenge of error accumulation, where irrelevant evidence may seep into the generation process, leading to incorrect responses Li et al.
-(2023a); Shi et al.
-(2023), a phenomenon referred to as external hallucination.
+A line of research harnesses the extensive parametric knowledge and robust logical reasoning capabilities of LLMs to ensure logical consistency either through self-reflection within a single model (Wang et al., 2023b; Ji et al., 2023b; Xu et al., 2023; Madaan et al., 2023; Dhuliawala et al., 2023) or through collaborative refinements or debates involving multiple models (Cohen et al., 2023; Du et al., 2023).
+Despite their strengths, LLMs are sometimes constrained by their knowledge boundaries and the complexity of the reasoning chain, resulting in occasional inaccuracies (Ren et al., 2023; Li et al., 2023b; Mallen et al., 2023) termed internal hallucination.
+To address this knowledge gap, retrieval-augmented generation methods leverage external knowledge as supplementary evidence to aid LLMs in providing accurate responses (Huang et al., 2023c; Ye et al., 2023; Niu et al., 2012; Xu et al., 2024; Varshney et al., 2023; Chern et al., 2023; Li et al., 2023a).
+However, these approaches, while effective, occasionally encounter the challenge of error accumulation, where irrelevant evidence may seep into the generation process, leading to incorrect responses (Li et al., 2023b; Shi et al., 2023), a phenomenon referred to as external hallucination.
 Our work only performs retrieval augmentation when hallucinations are detected, thereby maximizing the utilization of both the parametric knowledge and externally retrieved information.
-There are also some adaptive retrieval methods that assess the difficulty of questions or the confidence in responses to decide whether to retrieve documents Jiang et al.
-(2023); Mallen et al.
-(2023); Asai et al.
-(2023); Jeong et al.
-(2024).
-Due to space limitations, we discuss these works in Appendix C.
-3 Methodology
+3. Methodology
 Our objective is to enhance the factuality of LLM responses by integrating parametric and external knowledge.
@@
 For external hallucinations, Rowen resorts to external knowledge only when high uncertainties are found, ensuring that the final answer is both accurate and reliable.
-3.1 Stage 1: Generating Initial Answer
+| When responding to my question, please first evaluate the validity of the information or the assumption underlying the question.
+Once you’ve established its truth or existence, then proceed to deliver a detailed explanation or answer.
+Prioritize accuracy and fact-checking before diving into elaboration or conjecture.
+|
+3.1. Stage 1: Generating Initial Answer
 To maximize the exploitation of the parametric knowledge in LLMs, we initially employ their Chain-of-Thought (CoT) reasoning to generate a preliminary response.
-This process involves: critically assessing the validity of the information in the input query and prioritizing accuracy and fact-checking for answer elaboration, detailed in Appendix F Table 9.
-After generating the CoT thought , we ask the model to provide a concise answer for the input query .
-The answer will serve as the final response to the input query if our detection module ensures it is free from hallucinations.
-3.2 Stage 2: Deciding Whether to Retrieve
-To decide when to retrieve, we leverage model uncertainty, which refers to the confidence level of model outputs and serves as a crucial indicator for deciding when to trust LLMs Zhang et al.
-(2023d).
-Unfortunately, current consistency-based methods fail when LLMs provide consistent yet incorrect answers across different perturbations Zhang et al.
-(2023a).
+This process involves: critically assessing the validity of the information in the input query and prioritizing accuracy and fact-checking for answer elaboration, detailed in Table 1.
+After generating the CoT thought , we ask to provide a concise answer for the input query .
+Our aim is to ensure high-quality responses are generated from the outset.
+The answer is finalized as the ultimate response to the input query after our detection module ensures it is free from hallucinations.
+3.2. Stage 2: Deciding Whether to Retrieve
+To decide when to retrieve, we leverage model uncertainty, which refers to the confidence level of model outputs and serves as a crucial indicator for deciding when to trust LLMs (Zhang et al., 2023b).
+Unfortunately, current consistency-based methods fail when LLMs provide consistent yet incorrect answers across different perturbations (Zhang et al., 2023c).
 This issue may arise because these methods focus exclusively on semantic coherence within a single language or model.
@@
 If inconsistencies are detected in these responses, we flag them as potentially inaccurate and invoke a retrieval process.
-3.2.1 Diversifying Question Verbalizations
+3.2.1. Cross-Language / Model Perturbations
 To facilitate subsequent consistency-based hallucination detection, we begin by leveraging advancements in LLM prompting to generate semantically equivalent perturbations.
 Initially, we start with an input and instruct the model to provide a set of semantically equivalent questions .
-We use the prompt: "For the question [ORIGINAL QUESTION], please provide k semantically equivalent questions" with a high decoding temperature to generate diverse perturbed expressions.
+We use the prompt: ”For the question [ORIGINAL QUESTION], please provide k semantically equivalent questions” with a high decoding temperature to generate diverse perturbed expressions.
 After obtaining the diverse verbalized questions, we prompt the LM to generate its candidate responses according to the questions.
@@
 where is the length of the generated semantically equivalent questions .
-3.2.2 Cross-Language Consistency-based Detection
-Recent finding reveals that multilingual models do not always learn knowledge that is fully aligned across different languages Gao et al.
-(2024).
-Building on this insight, we propose that this inconsistency can be strategically used as an indicator for the necessity of information retrieval.
-When language models provide inconsistent answers to the same query in different languages, it signals a lack of certainty about the relevant knowledge.
-In such cases, retrieving external knowledge is crucial for accurate responses.
-To achieve cross-language detection, we first incorporate a cross-language consistency check (Rowen-CL) to evaluate the semantic consistency of responses to the same question across different languages.
-We introduce language-level perturbations by asking the model to translate the source-language questions into corresponding paraphrased questions in the target language.
+Cross-Language Detection
+To capture language-level uncertainty, we first incorporate a cross-language consistency check (Rowen-CL) to evaluate the semantic consistency of responses to the same question across different languages.
+To achieve this, we introduce language-level perturbations by asking the model to translate the source-language questions into corresponding paraphrased questions in the target language.
 The model is then instructed to generate corresponding answers to each question in the target language.
-To captures language-level cross-checking uncertainties, we utilize the generated questions and answers from all previous stages to calculate a numerical consistency score to calculate a numerical consistency score.
-Let denote a semantic equivalence checking operator that takes two QA pairs as inputs and returns "True" if they are semantically equivalent, and "False" otherwise.
+Cross-Model Detection
+Besides language-level cross-checking, we also introduce a cross-model detection module (Rowen-CM) to evaluate the semantic consistency of responses to the same question across different models.
+We adopt an additional verifier LM for model-level cross-checking and instruct the verifier LM to provide answers for each source question:
+3.2.2. Consistency Score Calculation
+In this step, we utilize the generated questions and answers from all previous stages to calculate a numerical consistency score that captures language-level and model-level cross-checking uncertainties.
+| Given the question Q, and two potential answers: answer A in English and answer B in Chinese.
+Your task is to determine if the content and meaning of A and B are equivalent in the context of answering Q.
+Consider linguistic nuances, cultural variations, and the overall conveyance of information.
+Respond with a binary classification.
+If A and B are equivalent, output ’True.’, otherwise output ’False’ |
+Cross-Language Consistency Score
+Let denote a semantic equivalence checking operator that takes two QA pairs as inputs and returns ”True” if they are semantically equivalent, and ”False” otherwise.
 We then map the response to a numerical semantic equivalent score: {"True" 1.0, "False" 0.0}.
-We leverage the model and utilize the prompt provided in Appendix F Table 10 to implement the cross-language checking operator and calculate the cross-language consistency score as:
+In our implementation, we leverage the model and utilize the prompt provided in Table 2 to implement the cross-language checking operator and calculate the cross-language consistency score as:
 where and denote the QA pairs in the source language and target language, respectively.
-3.2.3 Cross-Model Consistency-based Detection
-Besides language-level cross-checking, we also introduce a cross-model detection module (Rowen-CM) to evaluate the semantic consistency of responses to the same question across different models.
-We adopt an additional verifier LM for model-level cross-checking and instruct the verifier LM to provide answers for each semantically equivalent question:
-3.2.4 Hybrid Consistency-based Detection
+| Are the following two Question-Answer(QA) pairs semantically equivalent?
+Provide your best guess that it is correct (True or False).
+Given ONLY the guess (True or False), no other words or explanation.
+|
+Cross-Model Consistency Score
+Similar to the cross-language consistency score calculation, we use the prompt in Table 3 to implement the checking operator to calculate cross-model consistency score:
+where denote the QA pairs generated by verifier model .
+Hybrid Consistency Score
 The different variants of Rowen capture various aspects of uncertainty in the original response, complementing each other effectively.
@@
 where is a weight factor for the cross-model consistency score.
-3.3 Stage 3: Retrieval Augmented Generation
+3.3. Stage 3: Retrieval Augmented Generation
 If the consistency score falls below a threshold, it indicates possible hallucinated content in the original response . We then introduce a retrieval-augmented generation procedure.
@@
 To help the LM correct errors, we search for supporting evidence from external sources like online webpages.
-We first ask the model to generate search queries for each paraphrased question in , following the prompt in Appendix F Table 12.
+We first ask the model to generate search queries for each paraphrased question in .
 These queries are input into the online search engine to retrieve relevant knowledge, denoted as , used for correcting factual errors in .
@@
 With the retrieved evidence , the model reviews the original thought process and initial answer . The aim is to identify and correct inaccuracies, producing the refined answer :
-Appendix F Table 13 shows the prompt for repairing the original answer. The corrected answer serves as the final response to question .
-4 Experimental Setup
-4.1 Datasets and Evaluation Metrics
-TruthfulQA
-We use the TruthfulQA dataset Lin et al.
-(2022) to evaluate the ability of LLMs in generating truthful responses Zhang et al.
-(2023c); Kadavath et al.
-(2022).
-In our study, we focus on the generation task in TruthfulQA.
-Therefore, to evaluate the factuality of responses from LLMs, we calculate the GPT-Judge score, obtained by fine-tuning the babbage-002 model using the original fine-tuning data from their official repository222https://github.com/sylinrl/TruthfulQA.
-We also report the BLEU and Rouge-L scores to evaluate the lexical overlap between generated responses and ground-truth references.
-StrategyQA
-The StrategyQA dataset Geva et al.
-(2021) comprises crowdsourced yes / no questions that require multi-step reasoning for accurate answers.
-We follow previous work Jiang et al.
-(2023) to randomly sample 500 examples due to the cost consideration of running experiments.
-We also follow the settings of Wei et al.
-(2022) to generate both the reasoning process as well as the final answer.
-We present the exact-match accuracy of the generated yes / no answers compared to the gold-standard answers.
+The corrected answer serves as the final response to question .
 | Models | TruthfulQA | StrategyQA | ||
@@
 | Self-improvement Methods | ||||
-| CoVe Dhuliawala et al. (2023) | 48.01 | 12.81 | 26.52 | 61.40 |
-| Multi-agent Debate Du et al. (2023) | 50.83 | 3.94 | 21.05 | 65.73 |
-| Self-Reflection Ji et al. (2023b) | 42.99 | 3.86 | 18.18 | 62.40 |
+| CoVe (Dhuliawala et al., 2023) | 48.01 | 12.81 | 26.52 | 61.40 |
+| Multi-agent Debate (Du et al., 2023) | 50.83 | 3.94 | 21.05 | 65.73 |
+| Self-Reflection (Ji et al., 2023b) | 42.99 | 3.86 | 18.18 | 62.40 |
 | Retrieval-augmented Methods | ||||
-| Factool Chern et al. (2023) | 34.50 | 1.34 | 12.22 | 67.20 |
-| Detect-and-Mitigate Varshney et al. (2023) | 49.98 | 3.17 | 18.59 | 56.94 |
+| Factool (Chern et al., 2023) | 34.50 | 1.34 | 12.22 | 67.20 |
+| Detect-and-Mitigate (Varshney et al., 2023) | 49.98 | 3.17 | 18.59 | 56.94 |
 | Adaptive Retrieval Methods | ||||
-| FLARE Jiang et al. (2023) | 45.04 | 11.59 | 26.83 | 61.19 |
-| Adaptive-Retrieval Mallen et al. (2023) | 45.55 | 8.87 | 26.75 | 62.50 |
-| Self-RAG Asai et al. (2023) | 40.36 | 4.36 | 21.28 | 58.40 |
-| Adaptive-RAG Jeong et al. (2024) | 46.02 | 10.29 | 26.24 | 68.50 |
-| LUQ Zhang et al. (2024) | 55.08 | 5.79 | 21.44 | 71.00 |
+| FLARE (Jiang et al., 2023) | 45.04 | 11.59 | 26.83 | 61.19 |
+| Adaptive-Retrieval (Mallen et al., 2023) | 45.55 | 8.87 | 26.75 | 62.50 |
+| Self-RAG (Asai et al., 2023) | 40.36 | 4.36 | 21.28 | 58.40 |
+| Adaptive-RAG (Jeong et al., 2024) | 46.02 | 10.29 | 26.24 | 68.50 |
+| LUQ (Zhang et al., 2024) | 55.08 | 5.79 | 21.44 | 71.00 |
 | Our Framework | ||||
@@
 | Rowen-Hybrid | 59.34 | 15.27 | 31.15 | 75.60 |
-4.2 Baseline Methods
+4. Experimental Setup
+In this section, we outline the experimental setup of Rowen, detailing the datasets utilized, the evaluation metrics employed, the baseline methods considered, and the implementation specifics.
+4.1. Datasets and Evaluation Metrics
+We evaluate the hallucination mitigation performance on the following datasets:
+TruthfulQA
+We use the TruthfulQA dataset (Lin et al., 2022) to evaluate the ability of LLMs in generating truthful responses (Zhang et al., 2023a; Kadavath et al., 2022).
+In our study, we focus on the generation task in TruthfulQA.
+Therefore, to evaluate the factuality of responses from LLMs, we calculate the GPT-Judge score, obtained by fine-tuning the babbage-002 model using the original fine-tuning data from their official repository222https://github.com/sylinrl/TruthfulQA.
+We also report the BLEU and Rouge-L scores to evaluate the lexical overlap between generated responses and ground-truth references.
+StrategyQA
+The StrategyQA dataset (Geva et al., 2021) comprises crowdsourced yes / no questions that require multi-step reasoning for accurate answers.
+We follow previous work (Jiang et al., 2023) to randomly sample 500 examples due to the cost consideration of running experiments.
+We also follow the settings of Wei et al.
+(2022) to generate both the reasoning process as well as the final answer.
+We present the exact-match accuracy of the generated yes / no answers compared to the gold-standard answers.
+4.2. Baseline Methods
 We consider the following methods as our baselines: (1) Vanilla LLMs, such as ChatGPT.
-(2) Self-improvement methods: CoVe Dhuliawala et al.
-(2023) generates verification questions to self-analyze potential errors, systematically addressing each question to refine the baseline response.
-Self-Reflection Ji et al.
-(2023b) presents an interactive self-reflection methodology that incorporates knowledge acquisition and answer generation.
-Multi-agent Debate Du et al.
-(2023) utilize multiple LM agents to debate their individual responses over multiple rounds to arrive at a common final answer.
-(3) Retrieval-augmented methods: Factool Chern et al.
-(2023) leverages various tools to gather evidence about the factuality of the generated content.
-Detect and Mitigate Varshney et al.
-(2023) actively detects hallucinations during generation by identifying potential hallucination through the logit output values of LLMs.
-(4) Adaptive retrieval methods: FLARE Jiang et al.
-(2023) adopts an active retrieval strategy that only retrieves when LLMs generate low probability tokens.
-Adaptive-Retrieval Mallen et al.
-(2023) only retrieves when necessary based on a pre-defined threshold for entity popularity.
-Self-RAG Asai et al.
-(2023) trains a single arbitrary LM that adaptively retrieves passages on-demand.
-Adaptive-RAG Jeong et al.
-(2024) trains a query-complexity classifier to decide when to retrieve based on question complexity.
-LUQ Zhang et al.
-(2024) is a sampling-based uncertainty quantification for long text.
-4.3 Implementation Details
-In our experiments, we validate Rowen using the ChatGPT language model.
-We re-implement all baselines, except Self-RAG, using ChatGPT to ensure a fair comparison.
+(2) Self-improvement methods: CoVe (Dhuliawala et al., 2023) generates verification questions to self-analyze potential errors, systematically addressing each question to refine the baseline response.
+Self-Reflection (Ji et al., 2023b) presents an interactive self-reflection methodology that incorporates knowledge acquisition and answer generation.
+Multi-agent Debate (Du et al., 2023) utilize multiple LM agents to debate their individual responses over multiple rounds to arrive at a common final answer.
+(3) Retrieval-augmented methods: Factool (Chern et al., 2023) leverages various tools to gather evidence about the factuality of the generated content.
+Detect and Mitigate (Varshney et al., 2023) actively detects hallucinations during generation by identifying potential hallucination through the logit output values of LLMs.
+(4) Adaptive retrieval methods: FLARE (Jiang et al., 2023) adopts an active retrieval strategy that only retrieves when LLMs generate low probability tokens.
+Adaptive-Retrieval (Mallen et al., 2023) only retrieves when necessary based on a pre-defined threshold for entity popularity.
+Self-RAG (Asai et al., 2023) trains a single arbitrary LM that adaptively retrieves passages on-demand.
+Adaptive-RAG (Jeong et al., 2024) trains a query-complexity classifier to decide when to retrieve based on question complexity.
+LUQ (Zhang et al., 2024) is a sampling-based uncertainty quantification for long text.
+4.3. Implementation Details
+LLMs
+In our experiments, we validate Rowen using the GPT-3.5 language model, specifically the gpt-3.5-turbo-0613 version333https://api.openai.com/v1/chat/completions.
+We re-implement all baselines, except Self-RAG, using GPT-3.5 to ensure a fair comparison.
 For semantic perturbations, we configure the temperature to 1.0 to generate diverse perturbed expressions.
@@
 Considering the diversity of expressions and the latency in generating perturbations, we produce semantically equivalent questions.
-For the cross-language detection module, English serves as the source language while Chinese is employed as the target language.
-For the cross-model detection module, we adopt Qwen-Max-0428333https://qwenlm.github.io/blog/qwen-max-0428, a large instruction-tuned model for chat service, as the verifier LM.
-Following Chern et al.
-(2023), we utilize the Google Search API offered by Serper444https://serper.dev/ to search the top pages and extract the most pertinent search snippets from the API’s response.
-We also conduct extensive experiments to study the impact of various hyper-parameters in Appendix D.
-5 Experimental Results
-5.1 Main Results on Hallucination Mitigation
-We evaluate the effectiveness of Rowen on the TruthfulQA and StrategyQA datasets.
-Table 1 presents the overall performance of Rowen compared to several strong baselines.
-Rowen demonstrates superior performance on both datasets, with higher GPT-Judge score and accuracy, indicating the effectiveness of our proposed method.
+For the cross-language detection module, English serves as the source language while Chinese is employed as the target language444We study the impact of the choice of target language in § 5.4 Table 7..
+This decision is made considering the substantial cultural disparities between the two languages, which can enhance the model’s capability to detect semantic inconsistencies when responding to identical questions.
+For the cross-model detection module, we adopt Qwen-Max-0428555https://qwenlm.github.io/blog/qwen-max-0428, a large instruction-tuned model for chat service, as the verifier LM666We also study the impact of the choice of verifier LM in § 5.4 Table 8..
+We execute all experiments on 8 NVIDIA A800 80G GPUs.
 | Models | TruthfulQA | StrategyQA | ||||
@@
 | SAC3-all | 52.22 | 9.37 | 29.56 | 20.8 | 67.00 | 24.5 |
-| Rowen-Hybrid | 59.34 | 15.27 | 31.15 | 23.0 | 75.6 | 20.0 |
+| Rowen-Hybrid | 59.34 | 15.27 | 31.15 | 23.0 | 75.60 | 20.0 |
+Retrieval Module
+Following Chern et al.
+(2023), we utilize the Google Search API offered by Serper777https://serper.dev/ to search the top pages and extract the most pertinent search snippets from the API’s response.
+Subsequently, we parse the response to acquire various types of snippets, including answer boxes, knowledge graphs, and organic search results.
+5. Experimental Results
+In this section, we evaluate the effectiveness of Rowen for hallucination mitigation. In particular, we aim to answer the following research questions:
+RQ1: How does Rowen perform compared to existing strong methods for hallucination mitigation?
+RQ2: Is the proposed multilingual detection module superior to other hallucination detection methods?
+RQ3: How well does Rowen generalize to new datasets and models?
+RQ4: How do different hyperparameters affect the performance of Rowen?
+RQ5: Does Rowen eliminate both types of hallucinations as expected?
+RQ6: Does Rowen reduce unnecessary retrieval calls?
+5.1. Main Results (RQ1)
+We evaluate the effectiveness of Rowen on the TruthfulQA and StrategyQA datasets.
+Table 4 presents the overall performance of Rowen compared to several strong baselines.
+Rowen demonstrates superior performance on both datasets, with higher GPT-Judge score and accuracy, indicating the effectiveness of our proposed method.
 Vanilla ChatGPT shows a certain level of accuracy in answering factual questions, achieving scores of 47.92% and 61.40% on the two datasets, respectively.
@@
 These results underscore Rowen’s ability to effectively leverage parametric knowledge and external information for maximum advantage.
-5.2 Effect of Detection Module
-To validate the superiority of our hallucination detection module, we compare its performance in adaptive retrieval scenarios for hallucination mitigation with strong detection methods: (1) average token-level probability / entropy that utilizes the probabilities / entropies of tokens generated by a proxy LM (e.g., LLaMa2-7B) as a metric to measure hallucination.
-(2) SelfCheckGPT Manakul et al.
-(2023) that measures information consistency between the different responses to determine hallucinations.
-(3) Consistency-based method Zhang et al.
-(2023a), SAC3, that evaluates semantic-aware cross-check consistency, building upon the foundation of self-consistency principles.
-Based on the results in Table 2, it is evident that logits-based methods perform moderately well in detecting hallucinations, especially on the StrategyQA dataset.
+5.2. Effect of Detection Module (RQ2)
+To validate the effectiveness of our proposed hallucination detection module, we compare its performance in adaptive retrieval scenarios for hallucination mitigation with strong detection methods: (1) average token-level probability / entropy that utilizes the probabilities / entropies of tokens generated by a proxy LM (e.g., LLaMa2-7B) as a metric to measure hallucination.
+(2) SelfCheckGPT (Manakul et al., 2023) that measures information consistency between the different responses to determine hallucinations.
+(3) Consistency-based method (Zhang et al., 2023c), SAC3, that evaluates semantic-aware cross-check consistency, building upon the foundation of self-consistency principles.
+Based on the results in Table 5, it is evident that logits-based methods perform moderately well in detecting hallucinations.
 Specifically, the entropy-based estimation exhibits superior performance.
@@
 We also notice that Rowen achieves notable hallucination mitigation with a minimal number of retrieval calls.
-This underlines the superior efficiency of our hallucination detection module.
-5.3 Scalability of Rowen
+This underlines the superior efficiency of our adaptive retrieval module.
+5.3. Scalability of Rowen (RQ3)
 Scalability to Open-Source LLMs
-In addition to ChatGPT, we also assess Rowen’s effectiveness when employing open-source language models: Qwen1.5-14B-Chat, Qwen2-7B-Instruct, and Qwen2-72B-Instruct555https://huggingface.co/Qwen.
+In addition to ChatGPT, we also assess Rowen’s effectiveness when employing open-source language models: Qwen1.5-14B-Chat, Qwen2-7B-Instruct, and Qwen2-72B-Instruct888https://huggingface.co/Qwen.
 These models are chosen for the cross-language detection model due to their strong capabilities in following Chinese instructions, a critical feature for effective cross-language detection.
@@
 Instead, the Qwen-series models have demonstrated state-of-the-art performance in Chinese, which aligns with our research goals.
-For verifier LM , we choose to use Llama-3-8B-Instruct666https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct.
+For verifier LM , we choose to use Llama-3-8B-Instruct999https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct.
 Figure 3 shows the results of three variants of Rowen on three open-source models on TruthfulQA dataset.
@@
 | Rowen-Hybrid | 39.98 | 57.31 | 69.04 | 78.46 |
-5.4 Analysis of Computation Cost for Rowen
-Efficiency Analysis of LLM Calls
-To evaluate the efficiency of LLM calls in Rowen, we compare various methods—including Self-Reflection, Multi-agent Debate and three different Rowen variants—by analyzing the number of API calls required to answer a single question and their effectiveness in mitigating hallucinations.
-From the results shown in Table 4, we find that Rowen variants, particularly Rowen-CL and Rowen-CM, achieve significantly better hallucination mitigation compared to the baseline methods, while using a similar number of API call.
-To further improve efficiency, We design two useful speedup strategies to improve the efficiency of Rowen and describe them in Appendix E.
-| Methods | TruthfulQA | StrategyQA | ||
-| GPT-Judge | # Calls | Accuracy | # Calls | |
-| Self-Reflection | 42.99 | 6 | 62.40 | 5 |
-| Multi-agent Deb. | 50.83 | 6 | 65.73 | 6 |
-| Rowen-CL | 57.39 | 6 | 74.00 | 5 |
-| Rowen-CM | 56.29 | 5 | 72.40 | 4 |
-| Rowen-Hybrid | 59.34 | 8 | 75.60 | 6 |
-Efficiency Analysis of Retrieval Calls
-To verify the retrieval efficiency of Rowen, in Table 5, we compare the average number of retrieval calls made by Factool, Detect-and-Mitigate, FLARE, and Rowen-Hybrid to answer a question across two datasets.
-Rowen-Hybrid, which retrieves information only for uncertain responses, excels in both retrieval efficiency and factual accuracy, showcasing its superiority over other RAG methods.
-| Methods | # Num of Retrieval Calls | |
-| TruthfulQA | StrategyQA | |
-| Factool | 12.5 | 11.6 |
-| Detect-and-Mit. | 7.2 | 5.5 |
-| FLARE | 2.1 | 3.9 |
-| Adaptive-RAG | 0.9 | 1.4 |
-| Rowen-Hybrid | 1.5 | 0.5 |
-5.5 Quantitative Analysis of Hallucinations
-To verify whether Rowen effectively reduces internal and external hallucinations, we present a comparative analysis of the prevalence of both types of hallucinations across three methods—Factool, Detect-and-Mitigate, and Rowen-Hybrid.
-Internal hallucinations refer to the generation of wrong responses using only the parameterized knowledge of LLMs, while external hallucinations refer to the generation of incorrect responses using noisy documents introduced after retrieval.
-Figure 4 shows that Factool and Detect-and-Mitigate are significantly prone to both types of hallucinations.
-Both baseline methods have high levels of internal and external hallucinations, struggling with internal coherence and external fact alignment.
-In contrast, Rowen-Hybrid effectively reduces external hallucinations by timely integrating external knowledge, thereby avoiding unnecessary information retrieval and mitigating potential errors.
-6 Conclusion and Future Work
-In this paper, we introduce Rowen, a novel method designed to actively identify potential hallucinated responses in LLMs and correcting factual errors through an adaptive retrieval process.
-Our experiments on four datasets demonstrate the effectiveness of our approach.
-We also conduct detailed analytical experiments to underscore the impact of each module and the choices of hyper-parameters.
-Moving forward, we plan to expand our work by exploring how to effectively utilize retrieved evidence, even when irrelevant documents are integrated, and devise an effective method to minimize the influence of knowledge conflicts between internal knowledge and retrieved documents.
-Limitations
-Our method has several drawbacks.
-Firstly, it focuses on post-hoc correction, limiting its ability to detect hallucination only after generating a complete sentence rather than in real-time during the generation of contexts.
-Secondly, for ensuring accurate answers, our approach requires multiple requests to the API interface.
-This involves generating the CoT answer, perturbing and responding to the input, correcting the original answer, and so on, leading to increased latency.
-To address this, we can expedite the entire process by parallelizing API calls for both the OpenAI API and Google Search API.
-Additionally, reducing the number of perturbations can strike a better balance between detection precision and inference speed.
-Thirdly, our approach neglects to analyze and utilize the factual nature of the retrieved evidence, which we leave for future work.
-Ethics Statement
-This study delves into the detection and mitigation of hallucinations in language models (LMs), presenting broader implications for the realm of natural language processing (NLP) and addressing ethical concerns related to trust and reliability.
-Our method has the potential to enhance the accuracy and dependability of LMs, minimizing the risks associated with misinformation and biased outputs.
-Moreover, it foster accountability and trust in AI systems, contributing to the ongoing development of more reliable language models.
-Appendix A Kernel Density Estimation Visualization
-In our work, we hypothesize that lower consistency levels correlate with increased uncertainty in the model, consequently leading to a higher probability of inaccurate predictions.
-To verify this hypothesis, we analyze the consistency distribution through kernel density estimation plots on the TruthfulQA dataset.
-As illustrated in Figure 5, in the lower range of consistency check scores (left-skewed), the proportion of incorrect predictions is higher.
-This indicates that low consistency check scores are often associated with incorrect model predictions.
-On the contrary, in the higher range of consistency check scores (right-skewed), the proportion of correct predictions is higher.
-This indicates that high consistency check scores are usually associated with correct model predictions.
-There is some overlap between the distributions of different labels, but the overall trend shows that the higher the consistency check score, the higher the likelihood of correct predictions.
-These findings suggest that consistency check scores can be an effective indicator of prediction accuracy: the higher the score, the more accurate the model prediction is likely to be.
-These findings provide empirical support for our hypothesis.
-Additionally, we observe that the suitable threshold points for distinguishing correct from incorrect answers differ between Rowen-CL and Rowen-CM, with the optimal threshold for Rowen-CM being higher.
-This observation aligns with the conclusion in § D.
-Appendix B Related Work on Exploring Uncertainty for Hallucination Detection
-Uncertainty refers to the confidence level of the model outputs, and it serves an important indicator for identifying and eliminating hallucinations, so it can assist users in determining when to trust LLMs.
-In general, methods for exploring uncertainty for hallucination detection can be categorized into three types: (1) Logit-based estimation relies on accessing the model’s logits to calculate token-level probabilities or entropy, which are used to measure uncertainty Huang et al.
-(2023b); Varshney et al.
-(2023).
-However, this approach can pose challenges for black-box closed-source models.
-(2) Verbalized-based estimation involves prompting language models to express their uncertainty using specific prompts Xiong et al.
-(2023); Tian et al.
-(2023); Agrawal et al.
-(2023).
-However, these methods tend to display a high degree of overconfidence when expressing their confidence Xiong et al.
-(2023); Tao et al.
-(2024).
-(3) To overcome these limitations, consistency-based estimations are adopted to measure the consistency score among multiple responses provided by the model for a given question Manakul et al.
-(2023); Xiong et al.
-(2023); Wang et al.
-(2023b); Zhao et al.
-(2023).
-The underlying assumption suggests that when language models struggle with indecision and fabricate facts, they tend to provide logically inconsistent responses to identical questions.
-In this work, we propose that cross-language and cross-model consistency can offer highly sensitive signals for identifying hallucinations.
-Therefore, we utilize cross-language and cross-model detection modules that cross-check answers to the same question across different languages or models.
-This cross-checking paradigm serves as a powerful mechanism to identify hallucinations in LLMs.
-Appendix C Discussions
-There is a line of recent studies that concentrate on mitigating hallucinations with adaptive retrieval Jiang et al.
-(2023); Mallen et al.
-(2023); Asai et al.
-(2023); Jeong et al.
-(2024).
-Our work diverges from existing methodologies by: (1) Traditional self-consistency method Wang et al.
-(2023b) may falter when LLMs provide coherent yet incorrect responses to the same query, which is observed in SAC3.
-Our approach extends self-consistency uncertainty estimation to multilingual and multi-model settings, yielding superior hallucination mitigation performance compared to conventional methods (see Table 1).
-(2) FLARE Jiang et al.
-(2023) relies on the token probability of model outputs, a limitation for commercial closed-source models.
-In contrast, our method overcomes this constraint by assessing semantic similarity in generated content, which is suitable for white / black-box models.
-(3) Self-RAG Asai et al.
-(2023) necessitates intricate retraining processes and significant computational resources (4 Nvidia A100 GPUs and 80GB memory).
-In contrast, our method effectively mitigates hallucination without the need for retraining, underscoring its suitability for resource-constrained scenarios.
-(4) Adaptive-Retrieval Mallen et al.
-(2023) employs Wikipedia entity popularity for retrieval decisions, which proves ineffective for adversarial questions lacking clear entities in datasets like TruthfulQA.
-(5) Adaptive-RAG Jeong et al.
-(2024) dynamically decide whether to retrieve based on a pre-trained classifier, which is a smaller LM trained to predict the complexity level of incoming queries.
-Adaptive-RAG could be influenced by the accuracy of the classifier, given the absence of training datasets for query-complexity classifier, potentially leading to incorrect labeling of queries.
-Our approach, examining linguistic diversity (cross-language) and model response consistency (cross-model), effectively identifies and mitigates hallucinations.
-This enhances model robustness and reliability, providing a solid foundation for applying LLMs.
-Our method is novel, practical, flexible, training-free, and broadly applicable, offering new insights and optimizing LLM research and application.
-Appendix D Impact of Hyper-parameters in Rowen
-Impact of Detection Threshold.
+5.4. Impact of Hyper-parameters in Rowen (RQ4)
+We analyze the impact of key factors for hallucination mitigation on the TruthfulQA dataset, including detection thresholds, numbers of perturbed questions, and cross-model detection weights, as shown in Figure 4.
+Detection Threshold
 We experiment with various consistency thresholds to investigate their impact on hallucination mitigation effectiveness.
-As shown in Figure 6, we observe significant performance improvement when the detection threshold increases from 0.2 to 0.6, as this identifies more hallucinations and enhances factual accuracy by retrieving more external information.
+As shown in Figure 4(a), we observe significant performance improvement when the detection threshold increases from 0.2 to 0.6, as this identifies more hallucinations and enhances factual accuracy by retrieving more external information.
 However, further increasing the threshold introduces more noise and irrelevant evidence, degrading performance.
@@
 Additionally, Rowen-CL and Rowen-CM have different optimal thresholds of 0.6 and 0.8, respectively, and the Rowen-Hybrid method demonstrates the most stable performance across varying thresholds.
-Impact of Number of Perturbed Questions.
-We investigate the impact of the number of perturbed questions of Rowen on the performance of hallucination mitigation.
-The results in Figure 7 reveal that the mitigation performance improves with the increase in the number of perturbations, particularly notable in the transition from 4 to 6 perturbations.
-This enhancement is attributed to the greater number of perturbations aiding in higher detection accuracy.
-However, further increasing the perturbation count beyond 6 yields only marginal improvements in performance (even no improvements).
-Thus, after considering the results, we determine that the optimal number of perturbations for achieving the best hallucination mitigation performance is 6.
-Impact of Cross-Model Detection Weight.
+Number of Perturbed Questions
+We also study the effect of the number of perturbed question on hallucination mitigation and the results are shown in Figure 4(b).
+We observe that the performance of Rowen improves with an increasing number of question samples, yet the performance gain gradually diminishes after surpassing 6 question samples.
+This indicates that, in practice, using 6 question samples could achieve reasonably good performance at a relatively low computational cost.
+Cross-Model Detection Weight
 We also examine the impact of different cross-model detection weight for Rowen-Hybrid.
-We ensure that for any given , the detection threshold is recalibrated to maintain optimal performance.
-Figure 8 shows that GPT-Judge scores significantly increase as the weight rises from 0.25 to 1.0, indicating that introducing cross-model checking can improve the accuracy of hallucination detection.
-However, when the weight factor exceeds 1.0, performance declines, with scores dropping to around 58 at a weight of 1.5.
+Figure 4(c) shows that GPT-Judge scores significantly increase as the weight rises from 0.25 to 1.0, indicating that introducing cross-model checking can improve the accuracy of hallucination detection.
+However, when the weight factor exceeds 1.0, performance declines, with scores dropping to around 56 at a weight of 1.5.
 This suggests that while a moderate weight enhances factuality, excessively a high value allows the cross-model detection module to exert too much influence, reducing effectiveness.
@@
 | English - Chinese | High | 57.39 |
+| Verifier LM | Capability | GPT-Judge |
+| Qwen-Turbo | Low | 52.50 |
+| Qwen-Plus | Medium | 54.09 |
+| Qwen-Max | High | 56.29 |
 Impact of the Choice of Verifier LM
@@
 Specifically, Qwen-Max exhibited the highest effectiveness in this task, achieving an efficacy of 56.29%, while Qwen-Turbo and Qwen-Plus achieved 52.50% and 54.09%, respectively.
-These findings underscore the critical role of verifier LM selection in enhancing the effectiveness of hallucination mitigation.
-| Verifier LM | Capability | GPT-Judge |
-| Qwen-Turbo | Low | 52.50 |
-| Qwen-Plus | Medium | 54.09 |
-| Qwen-Max | High | 56.29 |
-Impact of Perturbation Objects
-In evaluating cross-language and cross-model perturbations, as well as calculating subsequent consistency scores, applying CL and CM perturbations to the original question appears to be a viable approach.
-However, recent studies suggest that merely perturbing questions might lead to responses that are consistent but not necessarily accurate Zhang et al.
-(2023a).
-This implies that question-level consistency alone does not suffice for assessing factual correctness, as these methods predominantly emphasize semantic coherence.
-To validate this, we conduct preliminary experiments by applying CL and CM perturbations to the original question (question-level consistency checking) and compare the results with those of Rowen.
-The results, as shown in Table 8, demonstrate that Rowen consistently outperforms the approach of applying perturbations to the original question.
-This reinforces that Rowen remains the most effective method for ensuring accuracy.
-| Object | Variants | TruthfulQA | StrategyQA |
-| GPT-Judge | Accuracy | ||
-| Input | CL | 55.34 | 72.00 |
-| CM | 55.85 | 72.00 | |
-| Perturbations | CL | 57.39 | 74.00 |
-| CM | 56.29 | 72.40 |
-Appendix E Strategies to improve efficiency
-We have implemented two helpful strategies to improve the efficiency without compromising the performance by:
-(1) Our cross-language detection and cross-model detection modules can operate in parallel.
-This significantly reduces latency, as it allows for the simultaneous evaluation of semantic consistency across different languages and models.
-(2) We have optimized the generation process for responding to multiple semantically equivalent perturbed questions by leveraging parallel API calls.
-This approach expedites the retrieval of diverse responses, which are crucial for our consistency checks.
-By employing these parallel processing strategies, we ensure that the efficiency of our method is improved while preserving its effectiveness.
-We believe these optimizations adequately address the concerns about the method’s efficiency and demonstrate our commitment to developing a scalable and high-performing framework.
-Appendix F Prompt Instruction Examples
-In this section, we provide these instructions used in our experiments.
-| When responding to my question, please first evaluate the validity of the information or the assumption underlying the question.
-Once you’ve established its truth or existence, then proceed to deliver a detailed explanation or answer.
-Prioritize accuracy and fact-checking before diving into elaboration or conjecture.
-|
-| Given the question Q, and two potential answers: answer A in English and answer B in Chinese.
-Your task is to determine if the content and meaning of A and B are equivalent in the context of answering Q.
-Consider linguistic nuances, cultural variations, and the overall conveyance of information.
-Respond with a binary classification.
-If A and B are equivalent, output ’True.’, otherwise output ’False’ |
-| Are the following two Question-Answer(QA) pairs semantically equivalent?
-Provide your best guess that it is correct (True or False).
-Given ONLY the guess (True or False), no other words or explanation.
-|
-| Transform the following question into a query suitable for information retrieval and then use the query to find relevant evidence. Follow these steps for a short but comprehensive approach: |
-| Original Question: |
-| 1. Simplify the question into key terms and concepts. |
-| 2. Remove any elements that are not essential for a search query. |
-| 3. If necessary, rephrase the question to focus on the most critical aspects for retrieval. |
-| You are an advanced AI with the ability to process and synthesize information efficiently. When provided with a question, you should consider the following aspects in your response: |
-| 1. The specific question asked: |
-| 2. The reasoning process: |
-| 3. The short answer previously given related to the question: |
-| 4. Evidences that have been retrieved: |
-| Use this information to generate a concise, accurate, and relevant answer that reflects the latest understanding of the topic based on the input provided. |
-| Your response should be clear, direct, and provide the most up-to-date information available while maintaining coherence with the previous discussion points. |
+These findings underscore the critical role of verifier LM selection in enhancing the effectiveness of hallucination mitigation tasks.
+5.5. Quantitative Analysis (RQ5)
+To verify whether Rowen effectively reduces internal and external hallucinations, we present a comparative analysis of the prevalence of both types of hallucinations across three methods—Factool, Detect-and-Mitigate, and Rowen-Hybrid.
+Internal hallucinations refer to the generation of wrong responses using only the parameterized knowledge of LLMs, while external hallucinations refer to the generation of incorrect responses using noisy documents introduced after retrieval.
+Figure 5 shows that Factool and Detect-and-Mitigate are significantly prone to both types of hallucinations.
+Both baseline methods have high levels of internal and external hallucinations, struggling with internal coherence and external fact alignment.
+In contrast, Rowen-Hybrid effectively reduces external hallucinations by timely integrating external knowledge, thereby avoiding unnecessary information retrieval and mitigating potential errors.
+| Methods | TruthfulQA | StrategyQA | ||
+| GPT-Judge | # Calls | Accuracy | # Calls | |
+| Self-Reflection | 42.99 | 6 | 62.40 | 5 |
+| Multi-agent Deb. | 50.83 | 6 | 65.73 | 6 |
+| Rowen-CL | 57.39 | 6 | 74.00 | 5 |
+| Rowen-CM | 56.29 | 5 | 72.40 | 4 |
+| Rowen-Hybrid | 59.34 | 8 | 75.60 | 6 |
+| Methods | # Num of Retrieval Calls | |
+| TruthfulQA | StrategyQA | |
+| Factool | 12.5 | 11.6 |
+| Detect-and-Mit. | 7.2 | 5.5 |
+| FLARE | 2.1 | 3.9 |
+| Adaptive-RAG | 0.9 | 1.4 |
+| Rowen-Hybrid | 1.5 | 0.5 |
+5.6. Analysis of Computation Cost (RQ6)
+Efficiency Analysis of LLM Calls
+To evaluate the efficiency of LLM calls in Rowen, we compare various methods—including Self-Reflection, Multi-agent Debate and three different Rowen variants—by analyzing the number of API calls required to answer a single question and their effectiveness in mitigating hallucinations.
+From the results shown in Table 9, we find that Rowen variants, particularly Rowen-CL and Rowen-CM, achieve significantly better hallucination mitigation compared to the baseline methods, while using a similar number of API call.
+Efficiency Analysis of Retrieval Calls
+To verify the retrieval efficiency of Rowen, in Table 10, we compare the average number of retrieval calls made by Factool, Detect-and-Mitigate, FLARE, and Rowen-Hybrid to answer a question across two datasets.
+Factool generates the most retrieval API calls due to verifying each claim.
+Detect-and-Mitigate and FLARE identify low-confident concepts in LLM outputs and call search APIs with fewer API calls.
+Adaptive-RAG underestimates the difficulty of adversarial questions in the TruthfulQA dataset, resulting in poor truthfulness scores despite using the minimum number of retrievals.
+Rowen-Hybrid, which retrieves information only for uncertain responses, excels in both retrieval efficiency and factual accuracy, showcasing its superiority over other RAG methods.
+This demonstrates the superiority of Rowen-Hybrid over other RAG methods in terms of both efficiency and effectiveness.
+6. Conclusion and Future Work
+In this paper, we introduce Rowen, a novel method designed to actively identify potential hallucinated responses in LLMs and correcting factual errors through an adaptive retrieval process.
+Our experiments on two datasets demonstrate the effectiveness of our approach.
+We also conduct detailed analytical experiments to underscore the impact of each module and the choices of hyper-parameters.
+Moving forward, we plan to expand our work by exploring how to effectively utilize retrieved evidence, even when irrelevant documents are integrated, and devise an effective method to minimize the influence of knowledge conflicts between internal knowledge and retrieved documents.
+Acknowledgement
+This work was supported by the Strategic Priority Research Program of the CAS under Grants No.XDB0680302, the National Natural Science Foundation of China (NSFC) under Grants No.
+62276248, the Key Research and Development Program of Xinjiang Uyghur Autonomous Region Grant No.
+2024B03026, the Beijing Nova Program under Grants No.
+20250484765, and the Youth Innovation Promotion Association CAS under Grants No.
+2023111.